# A Forward Reachability Perspective on Control Barrier Functions and Discount Factors in Reachability Analysis

### v3

@@ Abstract
Control invariant sets are crucial for various methods that aim to design safe control policies for systems whose state constraints must be satisfied over an indefinite time horizon. In this article, we explore the connections among reachability, control invariance, and Control Barrier Functions (CBFs). Unlike prior formulations based on backward reachability concepts, we establish a strong link between these three concepts by examining the inevitable Forward Reachable Tube (FRT), which is the set of states such that every trajectory reaching the FRT must have passed through a given initial set of states. First, our findings show that the inevitable FRT is precisely this initial set itself if it is a robust control invariant set with a differentiable boundary—a property necessary to connect with CBFs whose zero-level sets are control invariant. We highlight that if the boundary is not differentiable, the FRT of the robust control invariant set may become a strict superset of the invariant set and lose invariance. Next, we formulate a differential game between the control and disturbance, where the inevitable FRT is characterized by the zero-superlevel set of the value function. By incorporating a discount factor in the cost function of the game, the barrier constraint of the CBF naturally arises in the Hamilton-Jacobi equation and determines the optimal policy. Combining these results, the value function of our FRT formulation serves as a CBF-like function, and conversely, any valid CBF is also a forward reachability value function inside the control invariant set, thereby revealing the inverse optimality of the CBF. This strong link between reachability and barrier constraints is not achievable by previous backward reachability-based formulations, and addresses an important gap in existing literature for constructing valid CBFs to ensure safety.
@@
Table 1: Comparison of reachability methods discussed in the paper. More details are discussed in Section 5.
@@
1 Introduction
A typical way of characterizing a control invariant set is by using a scalar function whose zero-superlevel set defines the invariant set, known as the barrier certificate [34]. This concept has evolved into the notion of a control barrier function (CBF) [4], which mandates that the function satisfy a particular differential inequality condition. By enforcing this condition, control policies ensure not just the safety at the boundary of the set, but also a smooth deceleration of trajectories as they approach the boundary. We refer to this condition as the barrier constraint.
This paper delves into control invariance and the barrier constraint using a reachability approach, forging a strong link between reachability, control invariance, and CBFs. We will study systems with disturbance and robust control invariance, control invariance under the worst-case effect of the disturbance as described in [12, 38], and similarly the robust barrier constraint. Reachability analysis forms the basis for many methods that construct robust control invariant sets using dynamic programming principles [9, 28, 17, 2, 38, 13]. These methods compute the invariant sets by eliminating states that will inevitably reach unsafe regions, corresponding to backward reachability problems in which one is concerned with finding states that reach a set of terminal states. However, as we will present, these backward reachability methods only have weak linkage to the CBFs and the barrier constraint.
The key idea of our new formulation is to use the forward reachability concept, in which one is concerned with finding the states that a set of initial states reaches. Our core contribution is a discovery that the optimal policy of the forward reachability problem is determined by the barrier constraint, and that CBFs are forward reachability value functions.
To see this, first, we extend the notion of the minimal forward reachable tube (FRT) [27] to inevitable FRT, which encompasses states that are inevitably reached from the initial set, despite the worst-case disturbance. In this paper, we will use “FRT” to represent this inevitable FRT for brevity, unless specified otherwise. We then verify conditions under which the FRT remains identical to the initial set. We determine that the FRT remains unchanged when the initial set is 1) robustly control invariant, and 2) has a differentiable boundary. It is notable that if either these two conditions is not met, the FRT may differ from the initial set and not be robustly control invariant, whose examples are provided in this article. This result lays the foundation for interpreting CBFs as reachability value functions.
Towards this, we introduce a differential game that characterizes the inevitable FRT as the zero-superlevel set of the value function capturing the game between the control and disturbance. The value function is then determined as a viscosity solution [8] to a particular Hamilton-Jacobi partial differential equation (HJ-PDE) called the Hamilton-Jacobi forward reachable tube variational inequality (HJ-FRT-VI). The crux of this formulation is the incorporation of a discount factor in the cost function of the value function. As a result, first, the barrier constraint emerges in the HJ-FRT-VI, thus becoming the constraint that defines the optimal control policy of the FRT value function. Consequently, the value function acts as a CBF-like function in that it satisfies the barrier constraint almost everywhere. Conversely, any valid CBF also qualifies as a viscosity solution to the HJ-FRT-VI within the control invariant set and can therefore be interpreted as a forward reachability value function.
Additionally, the discount factor induces a contraction mapping in the Bellman operator of the value function, allowing the value function to be continuous and characterized as a unique viscosity solution to the HJ-FRT-VI. While the use of a discount factor in infinite-horizon optimal control problems is not novel [8], we emphasize that our key contribution lies in demonstrating how the discount factor leads to the emergence of the barrier constraint in the reachability formulation.
We highlight that prior formulations relying on backward reachability, introduced in [17, 2, 38, 13], were unable to establish such a connection. For instance, our earlier work in [13] incorporates the barrier constraint into a finite-horizon backward reachability problem, but its extension to the infinite-horizon setting fails to secure the continuity and boundedness of the value function. In contrast, our new formulation is the only method that not only secures the linkage to the barrier constraint but also ensures the boundedness and continuity of the value function and the solution uniqueness of the corresponding HJ-PDE (Table 1). Consequently, by adopting a forward reachability approach to control invariant sets in a manner not previously explored in the literature, our work creates a strong link between these three concepts and thus enables the use of Hamilton-Jacobi methods to characterize CBFs.
The rest of the article is organized as follows. In Section 2, we review the concepts of control invariance and CBFs, extending these notions to systems with disturbances. In Section 3, we introduce forward reachable tubes and present their application to robust control invariant sets. In Section 4, we detail the Hamilton-Jacobi formulation of the FRTs and establish a connection to CBFs. Section 5 compares our formulation and reachability formulations from prior work that have been applied to characterize control invariant sets. We conclude the article with closing remarks in Section 6.
[…]
2 Control Invariance and CBFs
@@ 2.1 Control Invariance
$\dot{\mathrm{x}}(t)=f(\mathrm{x}(t),\mathrm{u}(t))\;\text{for}\;t>0,\qquad\mathrm{x}(0)=x,$ (1)
where $x\in\mathbb{R}^{n}$ is an initial state, $\mathrm{x}:[0,\infty)\rightarrow\mathbb{R}^{n}$ is the solution to the ODE, and $\mathrm{u}:[0,\infty)\rightarrow U$ is a Lebesgue measurable control signal with $U\subset\mathbb{R}^{m_{u}}$. We use $\mathcal{U}$ to denote the set of Lebesgue measurable control signals. We assume that the control input set $U$ is compact, which holds for most physical systems whose actuation limit is bounded. Also, we assume that the system (1) satisfies the following conditions.
@@
    $f:\mathbb{R}^{n}\times U\rightarrow\mathbb{R}^{n}~{}\textnormal{is uniformly continuous,}$,
@@
By the above definition, a trajectory starting inside a control invariant set $S$ that is a subset of $X$ can remain within $S$ for all time, and therefore can stay safe in $X$. The control invariance of a set can be determined by a geometric relationship between the vector field and the tangent cone of the set, as defined next:
Definition 2 ((Bouligand’s) tangent cone [14]).
Given a closed set $S\subset\mathbb{R}^{n}$, the tangent cone to $S$ at $x\in S$ is defined as
$T_{S}(x):=\left\{z\in\mathbb{R}^{n}\mid\liminf_{\tau\rightarrow 0}\frac{\operatorname{dist}(x+\tau z,S)}{\tau}=0\right\},$ (2)
where $\operatorname{dist}(y,S):=\min_{z\in S}{\left\lVert y-z\right\rVert}$.
The tangent cone captures the feasible directions in which one can move from the point $x$ within the set $S$.
@@
(Tangential characterization of closed control invariant sets [5, Theorem 11.3.4]) Let the dynamics (1) satisfy Assumption 1. Then, a closed set $S\subset\mathbb{R}^{n}$ is (forward) control invariant under the dynamics (1) if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{such that}~{}f(x,u)\in T_{S}(x).$ (3)
Below is a corollary of the lemma in a special case when the set $S$ has a differentiable boundary (Assumption 2), by introducing a scalar function $h_{S}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies Assumption 3.
Assumption 2.
$S$ is a closed set whose interior is non-empty, and whose boundary, $\partial S$, is continuously differentiable. [^1] For each point $x\in\partial S$, there exists $r>0$ and a $C^{1}$ function $\eta:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ such that $S\cap B_{r}(x)=\{x\in B_{r}(x)~{}|~{}x_{n}\geq\eta(x_{1},...,x_{n-1})\}$, where relabeling and reorienting the coordinates axes are allowed [15].
Assumption 3.
Given a closed set $S$, $h_{S}:\mathbb{R}^{n}\to\mathbb{R}$ is a function whose zero-superlevel set is $S$, i.e. $S=\{x\in\mathbb{R}^{n}~{}|~{}h_{S}(x)\geq 0\}$, and satisfies the following conditions:
@@
    $\mathrm{Int}(S)$ =$=\{x\in\mathbb{R}^{n}~{}|~{}h_{S}(x)>0\},$
    $\partial S$ =$=\{x\in\mathbb{R}^{n}~{}|~{}h_{S}(x)=0\}.\vspace{-1em}$ (4)
@@
    (Differentiability and boundedness) $h_{S}$ is uniformly continuously differentiable and both upper and lower bounded.
@@
    $\frac{\partial h_{S}}{\partial x}(x)\neq 0\quad\forall x\in\partial S+B_{\varepsilon}.\vspace{-0.5em}$ (5)
@@
Under Assumptions 1, 2, and 3, $S$ is (forward) control invariant under the dynamics (1) if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{such that}~{}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u)\geq 0.\vspace{-0.5em}$ (6)
@@
$T_{S}(x)=\left\{z\in\mathbb{R}^{n}|\frac{\partial h_{S}}{\partial x}\cdot z\geq 0\right\},\vspace{-0.5em}$ (7)
when Assumptions 2 and 3 hold. ∎
Lemma 2 is known as Nagumo’s theorem for autonomous systems [31]. For a given $S$ that is control invariant, (6) holds for any $h_{S}$ satisfying Assumption 3. The specific choice of $h_{S}$ does not affect condition (6). Also, an $h_{S}$ satisfying Assumption 3 always exists for the set $S$ satisfying Assumption 2, by selecting a regularized distance function for $S$ [25, Theorem 2.1].
Note that a control invariant set does not necessarily have a differentiable boundary. In general, the maximal control invariant set contained in the desired safety constraint set $X$ might have a non-differentiable boundary [5]. However, the differentiability of the boundary will render a few noticeable differences in the theory that will be developed in Section 3, necessitating Assumption 2 for our main theorems.
Next, we consider the concept of control invariance extended to systems with disturbance. There exist various formulations of robustness with respect to disturbances or uncertainties in system dynamics [10, 21, 22, 24]. In this paper, we employ the differential game-based formulation that interprets the disturbance as an adversarial agent playing against the control input [16], as commonly done in the Hamilton-Jacobi analysis for systems with bounded disturbance [12, 17, 38].
For this, we consider the system dynamics
$\dot{\mathrm{x}}(t)=f(\mathrm{x}(t),\mathrm{u}(t),\mathrm{d}(t))\;\text{for}\;t>0,\qquad\mathrm{x}(0)=x,\vspace{-0.5em}$ (8)
where $\mathrm{d}:[0,\infty)\rightarrow D$ is a Lebesgue measurable disturbance signal and $D\subset\mathbb{R}^{m_{d}}$ is a compact set. Note that control systems (1) can be regarded as a special case when the disturbance set $D$ in (8) is set to a singleton (e.g. $D=\{0\}$). We use $\mathcal{D}$ to denote the set of Lebesgue measurable disturbance signals. We assume conditions on the dynamics, similar to Assumption 1:
@@
    $f:\mathbb{R}^{n}\times U\times D\rightarrow\mathbb{R}^{n}~{}\textnormal{is uniformly continuous,}$,
@@
$\begin{split}\Xi_{d}\coloneqq\{&\xi_{d}:\mathcal{U}\rightarrow\mathcal{D}~{}|~{}\forall s\in[0,\infty)\text{ and }\mathrm{u},\bar{\mathrm{u}}\in\mathcal{U},\\ &\text{if }\mathrm{u}(\tau)=\bar{\mathrm{u}}(\tau)\text{ a.e. }\tau\in[0,s],\;\\ &\text{then }\xi_{d}[\mathrm{u}](\tau)=\xi_{d}[\bar{\mathrm{u}}](\tau)\text{ a.e. }\tau\in[0,s]\}.\end{split}\vspace{-0.5em}$ (9)
@@
Similar to Lemma 2, robustly control invariant sets can be verified by examining the vector field of the dynamics at the boundary of the sets:
@@
(Tangential characterization of robustly control invariant sets) Under Assumptions 2, 3, and 4, $S$ is robustly (forward) control invariant under the dynamics (8) if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{such that}~{}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)\geq 0\;\forall d\in D.$ (10)

### v4

@@ Abstract
Control invariant sets are crucial for various methods that aim to design safe control policies for systems whose state constraints must be satisfied over an indefinite time horizon. In this article, we explore the connections among reachability, control invariance, and Control Barrier Functions (CBFs). Unlike prior formulations based on backward reachability concepts, we establish a strong link between these three concepts by examining the inevitable Forward Reachable Tube (FRT), which is the set of states such that every trajectory reaching the FRT must have passed through a given initial set of states. First, our findings show that the inevitable FRT is a robust control invariant set if it has a continuously differentiable boundary. If the boundary is not differentiable, the FRT may lose invariance. We also show that any robust control invariant set including the initial set is a superset of the FRT if the boundary of the invariant set is differentiable. Next, we formulate a differential game between the control and disturbance, where the inevitable FRT is characterized by the zero-superlevel set of the value function. By incorporating a discount factor in the cost function of the game, the barrier constraint of the CBF naturally arises in the Hamilton-Jacobi (HJ) equation and determines the optimal policy. The resulting FRT value function serves as a CBF-like function, and conversely, any valid CBF is also a forward reachability value function. We further prove that any $C^{1}$ supersolution of the HJ equation for the FRT value functions is a valid CBF and characterizes a robust control invariant set that outer-approximates the FRT. Building on this property, finally, we devise a novel method that learns neural control barrier functions, which learn an control invariant superset of the FRT of a given initial set.
@@
1 Introduction
Control invariant sets can be computed using set-theoretic and optimization-based approaches. For example, convex and scalable methods for computing robust control invariant sets of nonlinear systems have been proposed in [20, 41, 15, 10]. These approaches offer computational advantages and scalability, but typically rely on restrictive assumptions such as convexity, or specific set parameterizations.
In contrast, approaches that characterize control invariant sets using a scalar function—referred to as the barrier certificate [39], do not explicitly require such assumptions. This concept has evolved into the notion of a control barrier function (CBF) [4], which mandates that the function satisfy a particular differential inequality condition. By enforcing this condition, control policies ensure not just the safety at the boundary of the set, but also a smooth deceleration of trajectories as they approach the boundary. We refer to this condition as the barrier constraint. Although these concepts provide an effective framework for safety verification and control, designing CBFs remains a nontrivial problem, as the feasibility of the barrier constraint must be ensured everywhere inside the safe set. This challenge has motivated significant recent efforts toward systematic methods for designing or learning CBFs [17].
This paper focuses on reachability analysis, which is closely related to robust CBFs, namely control invariance under adversarial disturbances [12, 49]. As summarized in Table 1, many reachability-based methods for robust control invariance rely on backward reachability [33, 21, 2, 49, 50, 13], which characterizes the set of initial states that can remain in a given set despite disturbances. However, these approaches do not directly yield CBFs. Enforcing the barrier condition modifies the associated Hamilton–Jacobi partial differential equation (HJ-PDE), which may result in unbounded or discontinuous value functions and multiple solutions. Consequently, numerical solutions may not necessarily correspond to the robust control invariant set.
Unlike prior reachability methods, our approach leverages forward reachability, which characterizes the states reachable from a given set of initial conditions. The key insight is that this perspective leads to a formulation whose value function naturally satisfies the requirements of control barrier functions, remaining bounded and continuous while enforcing the barrier condition.
To formalize this idea, we extend the notion of the minimal forward reachable tube (FRT) [34] to the inevitable FRT, which represents the set of states that are inevitably reached from the initial set despite worst-case disturbances. For brevity, we use the term “FRT” to denote this inevitable FRT unless specified otherwise. The FRT is characterized through an HJ formulation, which we call the HJ forward reachable tube variational inequality (HJ-FRT-VI).
Next, we show that the FRT is robustly control invariant if its boundary is continuously differentiable. Otherwise, robust control invariance may not hold because the barrier constraint is satisfied almost everywhere on the boundary but not everywhere. Moreover, the viscosity solution of the proposed HJ-FRT-VI may contain non-differentiable points and therefore cannot directly serve as a CBF, even though it satisfies the barrier condition almost everywhere in the state space.
To address this issue, we prove that any $C^{1}$ continuous supersolution of the proposed HJ-FRT-VI is a valid CBF and characterizes a robust control invariant set that outer-approximates the FRT. Based on this result, we further develop a learning-based numerical method to compute such supersolutions as CBFs.
We highlight the following two main theoretical contributions.
  1. 1.
    This paper establishes that CBFs can be systematically constructed using forward reachability, in contrast to existing backward-reachability-based approaches [33, 21, 2, 49, 13] (see Table 1). The proposed formulation yields a value function that satisfies the barrier constraint while remaining bounded, continuous, and uniquely defined as the solution of the proposed HJ-FRT-VI.
  2. 2.
    This paper establishes a novel connection between supersolutions of the HJ-FRT-VI and control barrier functions. While supersolutions are a classical concept in the HJ-PDE literature [19], their control-theoretic interpretation in the context of safety and control invariance has not been previously characterized. This paper proves that any $C^{1}$ supersolution of the proposed HJ-FRT-VI is a valid CBF and characterizes a robust control invariant set that outer-approximates the FRT. This result assigns a concrete safety and invariance meaning to supersolutions and enables the systematic construction of differentiable CBFs.
The rest of the article is organized as follows. In Section 2, we review the concepts of control invariance, CBFs, and backward reachability analysis for systems with disturbances. In Section 3, we introduce forward reachable tubes and present their relevance to robust control invariance. In Section 4, we detail the Hamilton-Jacobi formulation of the FRTs and establish a connection to CBFs. In Section 5, we propose a method for learning CBFs developed upon the theoretical properties in Sections 3 and 4. Section 6 provides a discussion on how our work is distinguished from the reachability formulations in prior work. We conclude the article with closing remarks in Section 7.
Notation: $\left\lVert\cdot\right\rVert$ indicates the $l^{2}$ norm in the Euclidean space. For two same dimensional vectors $a$ and $b$, $a\cdot b$ denotes the inner product. For a set $A$, $\mathrm{Int}(A)$, $\partial A$, $A^{c}$ denote the interior, the boundary, and the complement of $A$, respectively. For a point $x\in\mathbb{R}^{n}$ and $r>0$, we define $B_{r}(x)$ as the hyperball centered at $x$ with radius $r$, $B_{r}(x):=\{y\in\mathbb{R}^{n}\;|\;\left\lVert y-x\right\rVert\leq r\}$. For $\varepsilon>0$ and a set $A$, $A+B_{\varepsilon}\coloneqq\bigcup_{x\in A}B_{\varepsilon}(x)$, and $A-B_{\varepsilon}\coloneqq A\setminus\bigcup_{x\in A^{c}}B_{\varepsilon}(x)$.
@@ 1 Introduction
Table 1: Comparison of reachability methods discussed in the paper. Background in Section 2 and discussion in Section 6.
@@
2 Background
@@ 2.1 Control Invariance
$\dot{\mathrm{x}}(t)=f(\mathrm{x}(t),\mathrm{u}(t))\;\text{for}\;t>0,\qquad\mathrm{x}(0)=x,\vskip-5.0pt$ (1)
where $x\in\mathbb{R}^{n}$ is an initial state, $\mathrm{x}:[0,\infty)\rightarrow\mathbb{R}^{n}$ is the solution to the ODE, and $\mathrm{u}:[0,\infty)\rightarrow U$ is a Lebesgue measurable control signal with $U\subset\mathbb{R}^{m_{u}}$. We assume that $U$ is compact and use $\mathcal{U}$ to denote the set of Lebesgue measurable control signals. We also assume that the system (1) satisfies the following conditions.
@@
    $f:\mathbb{R}^{n}\times U\rightarrow\mathbb{R}^{n}~\textnormal{is uniformly continuous,}$
@@
By the above definition, a trajectory starting inside a control invariant set $S$ that is a subset of $X$ can remain within $S$ for all time, and therefore can stay safe in $X$. The control invariance can be determined by a geometric relationship between the vector field and the tangent cone of the set:
@@
(Tangential characterization of closed control invariant sets [5, Theorem 11.3.4]) A closed set $S\subset\mathbb{R}^{n}$ is (forward) control invariant under the dynamics (1) if and only if for all $x\in\partial S$,
$\exists u\in U~\textnormal{such that}~f(x,u)\in T_{S}(x),\vskip-5.0pt$ (2)
where $T_{S}(x)$ is the tangent cone to SS. [^1] Given a closed set $C\subset\mathbb{R}^{n}$, the (Bouligand’s) tangent cone to $C$ at $x\in C$ is defined as $T_{C}(x):=\left\{z\in\mathbb{R}^{n}\mid\liminf_{\tau\rightarrow 0}\frac{\operatorname{dist}(x+\tau z,C)}{\tau}=0\right\},\vskip-5.0pt$ (3) where $\operatorname{dist}(y,C):=\min_{z\in C}{\left\lVert y-z\right\rVert}$ [14]. The tangent cone captures the feasible directions in which one can move from the point $x$ within the set $C$.
We consider a corollary of the lemma in a special case when the set $S$ has a differentiable boundary (Condition 1), by introducing a scalar function $h_{S}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies Condition 2.
Condition 1 (Differentiability of the boundary).
$C$ is a closed set whose interior is non-empty, and whose boundary, $\partial C$, is continuously differentiable. [^2] For each point $x\in\partial C$, there exists $r>0$ and a $C^{1}$ function $\eta:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ such that $C\cap B_{r}(x)=\{x\in B_{r}(x)~|~x_{n}\geq\eta(x_{1},...,x_{n-1})\}$, where relabeling and reorienting the coordinates axes are allowed [19].
Condition 2.
Given a closed set $C$, $h_{C}:\mathbb{R}^{n}\to\mathbb{R}$ is a function whose zero-superlevel set is $C$, i.e. $C=\{x\in\mathbb{R}^{n}~|~h_{C}(x)\geq 0\}$, and satisfies the following conditions:
@@
    $\mathrm{Int}(C)$ $=\{x\in\mathbb{R}^{n}~|~h_{C}(x)>0\},$
    $\partial C$ $=\{x\in\mathbb{R}^{n}~|~h_{C}(x)=0\}.\vskip-10.00002pt$ (4)
@@
    (Differentiability and boundedness) $h_{C}$ is uniformly continuously differentiable and both upper and lower bounded.
@@
    $\frac{\partial h_{C}}{\partial x}(x)\neq 0\quad\forall x\in\partial C+B_{\varepsilon}.\vskip-5.0pt$ (5)
Proposition 1.
For any $h_{C}$ satisfying Condition 2, its zero-superlevel set $C$ satisfies Condition 1 [40, Theorem 9.28], and for any set $C$ satisfying Condition 1, $h_{C}$ satisfying Condition 2 always exists [30, Theorem 2.1].
Given these conditions, Lemma 1 results in the following corollary:
@@
$S$ satisfying Condition 1 is (forward) control invariant under the dynamics (1) if and only if for all $x\in\partial S$,
$\exists u\in U~\textnormal{such that}~\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u)\geq 0,\vskip-5.0pt$ (6)
for $h_{S}$ satisfying Condition 2.
@@
$T_{S}(x)=\left\{z\in\mathbb{R}^{n}|\frac{\partial h_{S}}{\partial x}\cdot z\geq 0\right\}.\vskip-5.0pt$ (7)
Lemma 2 is known as Nagumo’s theorem for autonomous systems [36]. For a given $S$ that is control invariant, (6) holds for any $h_{S}$ satisfying Condition 2. The specific choice of $h_{S}$ does not affect condition (6).
Next, we extend the concept to systems with disturbance. There exist various formulations of robustness with respect to disturbances or uncertainties in dynamics [9, 25, 27, 29]. In this paper, we employ the differential game that interprets the disturbance as an adversary playing against the control input [18], as commonly done in the Hamilton-Jacobi analysis [12, 21, 49]. For this, we consider the system dynamics
$\dot{\mathrm{x}}(t)=f(\mathrm{x}(t),\mathrm{u}(t),\mathrm{d}(t))\;\text{for}\;t>0,\qquad\mathrm{x}(0)=x,\vskip-5.0pt$ (8)
where $\mathrm{d}:[0,\infty)\rightarrow D$ is a Lebesgue measurable disturbance signal and $D\subset\mathbb{R}^{m_{d}}$ is a compact set. Control systems (1) can be regarded as a special case when the disturbance set $D$ in (8) is set to a singleton (e.g. $D=\{0\}$). We use $\mathcal{D}$ to denote the set of Lebesgue measurable disturbance signals. We assume conditions on the dynamics, similar to Assumption 1:
@@
    $f:\mathbb{R}^{n}\times U\times D\rightarrow\mathbb{R}^{n}~\textnormal{is uniformly continuous,}$
@@
[…]
@@
Similar to Lemma 2, robust control invariant sets can be verified by examining the vector field of the dynamics at the boundary of the sets:
@@
(Tangential characterization of robust control invariant sets) $S$ satisfying Condition 1 is robustly (forward) control invariant under the dynamics (8) if and only if for all $x\in\partial S$,
$\exists u\in U~\textnormal{such that}~\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)\geq 0\;\forall d\in D,\vskip-5.0pt$ (10)
for $h_{S}$ satisfying Condition 2.